# New Characteristics and Modeling of 6G Channels: A Unified Channel Model towards Standardization

### v1

@@ Abstract
As 6G research advances, the growing demand leads to the emergence of novel technologies such as Integrated Sensing and Communication (ISAC), new antenna arrays like Extremely Large MIMO (XL-MIMO) and Reconfigurable Intelligent Surfaces (RIS), along with multi-frequency bands (6-24 GHz, above 100 GHz). Standardized unified channel models are crucial for research and performance evaluation across generations of mobile communication, but the existing 5G 3GPP channel model based on geometry-based stochastic model (GBSM) requires further extension to accommodate these 6G technologies. In response to this need, this article first investigates six distinctive channel characteristics introduced by 6G techenologies, such as ISAC target RCS, sparsity in the new mid-band, and others. Subsequently, an extended GBSM (E-GBSM) is proposed, integrating these characteristics into a unified modeling framework. The proposed model not only accommodates 6G technologies with flexibility but also maintains backward compatibility with 5G, ensuring a smooth evolution between generations. Finally, the implementation process of the proposed model is detailed, with experiments and simulations validate its effectiveness and accuracy, providing support for 6G channel modeling standardization efforts.
@@ I Introduction
The sixth generation (6G) of wireless communication technology aspires to redefine connectivity, offering transformative advancements such as higher data rates, ultra-low latency, and ubiquitous coverage [1]. Given the complexity of 6G systems, standardization serves as a cornerstone for bridging diverse technological advancements, and global standardization plays a pivotal role in ensuring compatibility, interoperability, and scalability across diverse systems and regions. Notably, since 2020, the International Telecommunication Union (ITU) has spearheaded 6G research, targeting deployment by 2030, and finalized its 6G Development Framework in 2023 [2]. Concurrently, the World Radiocommunication Conference (WRC) is addressing spectrum requirements, with allocations expected by 2027. The 3rd Generation Partnership Project (3GPP) has also launched Release 19 to initiate 6G standardization efforts. As with previous generations, channel modeling remains foundational, providing the models essential for uniform benchmarking, system design, and performance evaluation.
The evolution of channel models from 1G to 6G is summarized in Table I, highlighting the transition from empirical models to standardized approaches. For 2G and earlier generations, empirical path loss models such as the Okumura-Hata model and the COST-231 model were predominant. Starting from 3G, standardized channel models have been introduced for each generation of mobile communication. The 3G era saw the introduction of the Spatial Channel Model (SCM), combining statistical modeling with spatial correlation to support frequency-selective fading. With the widespread adoption of MIMO technology, 4G introduced the SCME and WINNER models, which further enhanced spatial modeling capabilities. For 5G, the Geometry-Based Stochastic Model (GBSM) integrated geometric and stochastic properties[3], while the 3D-GBSM extended channel representation into the vertical domain [4].
TABLE I: Development history of channel models from 1G to 6G
@@
Dimension Time Dom. Time-frequency Dom. Time-frequency Dom.
@@
6G aims to transform communication systems by expanding application scenarios and adopting advanced technologies to meet unprecedented performance demands. The IMT-2030 Framework [2] outlines six 6G use cases, building on three primary 5G scenarios, with 15 performance indicators surpassing 5G in sensing, positioning, data rates, and coverage. Several novel technologies have been proposed to achieve these goals. Integrated Sensing and Communication (ISAC) combines sensing and communication functions for precise positioning and autonomous driving. Advanced antenna technologies, such as Extremely Large MIMO (XL-MIMO) and Reconfigurable Intelligent Surfaces (RIS), enhance spectral efficiency, capacity, and reliability. XL-MIMO utilizes extensive antenna arrays, while RIS actively manipulates electromagnetic waves to improve coverage. To support these capabilities, new frequency ranges, such as the 6-24 GHz, provide the bandwidth and spectrum resources required for achieving 6G’s peak data rates [5].
While the 5G channel model, grounded in GBSM, has successfully supported key technologies like MIMO and millimeter-wave communication, it now cannot be directly applied to 6G due to the expanded technical requirements, broader frequency ranges, and more diverse scenarios introduced by 6G technologies. ISAC systems, for instance, necessitate target modeling to capture the characteristics of both communication and sensing channels[6]. XL-MIMO’s large antenna arrays introduce spatial non-stationarity (SnS) and near-field effects[7], significantly altering channel behavior. RIS-assisted communications present challenges in modeling the Transmitter(Tx)-RIS and RIS-Receiver(Rx) channel coupling, while new frequency bands, such as mid-band [8], millimeter-wave, and sub-THz, exhibit sparse multipath propagation[9].
@@
In response to these challenges, this article investigates the channel characteristics introduced by key 6G technologies and, drawing on discussions from the 3GPP Technical Specification Group Radio Access Network Working Group 1 (TSG RAN1), proposes a unified channel modeling framework based on an extended geometry-based stochastic model (E-GBSM). This framework integrates the distinctive features of ISAC, advanced antenna technologies, and new frequency bands (new mid-band and above 100 GHz), while ensuring compatibility with existing standards. By extending current models, it maintains backward compatibility, provides a clear implementation roadmap, and addresses critical requirements such as wideband support and multi-technology evaluation.
@@ II-A ISAC
Compared to traditional communication channels, ISAC channels emphasize the behavior and characteristics of sensing targets. Measurement studies focus on three aspects: (1) RCS properties of sensing targets, (2) modeling of the target channel and its sub-channels (Tx-Target and Target-Rx), and (3) correlations between sensing target channels and communication channels.
The first aspect involves the investigation of RCS properties. Although extensively studied in radar systems, ISAC targets differ in frequency bands and scenarios, necessitating further research [11]. Typical targets include UAVs, humans, and vehicles, relevant to applications such as UAV detection and autonomous driving assistance. Measurements reveal distinct RCS distributions, e.g., UAVs and humans exhibit weak angular dependence, while vehicles show four peaks corresponding to their sides. RCS also varies with frequency and distance, stabilizing at longer ranges. Certain studies highlight potential correlations between bistatic and monostatic RCS within specific angular ranges, simplifying modeling.
@@ II-B New antenna
The channel characteristics of advanced antenna technologies, such as XL-MIMO and RIS, focus on near-field effects, spatial non-stationarity (SnS), and multi-segment concatenated channels in RIS systems.
@@ II-C New frequency
As frequency increases, signal wavelengths shorten, reducing their ability to reflect off or penetrate surfaces, which leads to a higher proportion of scattered paths with weaker energy compared to direct paths. This results in received signal power becoming concentrated in the direct path and a few dominant non-line-of-sight (NLOS) paths, creating sparse channel characteristics. Various metrics, such as channel rank or spatial degrees of freedom, quantify sparsity. This article adopts the Gini Index, as proposed in [9], to measure sparsity, where a Gini Index closer to 1 (within the 0-1 range) indicates greater sparsity. Channel measurements from [9] across centimeter-wave, new mid-band, and THz bands show that the Gini Index increases with frequency, confirming that higher frequencies exhibit more pronounced sparsity. This sparsity appears at both the cluster and path levels: at higher frequencies, lower-energy scattering clusters become harder to detect, leading to fewer observable clusters, while at the path level, power concentrates within a single dominant path in each cluster. These findings indicate that existing 3GPP models, which assume equal power distribution across paths, are inadequate for new frequencies.
Figure 2: The illustration of the unified channel model for 6G.
@@ III-A Consideration for 6G channel model
Support for evaluating multiple technologies. While the 5G channel model accommodates several features like spatial consistency, temporal evolution, and blockage effects (e.g., V2X networks and oxygen absorption for 53-67 GHz), it lacks provisions for many 6G technologies. For instance, ISAC requires models that capture sensing channels from the Tx to the target and then to the Rx, which are unsupported in current models like 3GPP TR 38.901 [3]. Similarly, RIS models must represent concatenated Tx-RIS-Rx channels, and XL-MIMO demands adjustments to incorporate non-stationary and near-field effects, which traditional models based on spatial stationarity and far-field assumptions fail to address.
Support for a wide frequency range. Achieving 6G’s peak data rates requires both higher frequencies (e.g., THz bands [12]) and better utilization of underused bands (e.g., 6-24 GHz). Current models extrapolate 6-24 GHz parameters from FR1 and FR2 band [13] and lack representation of the frequencies above 100 GHz. Moreover, they do not account for frequency-dependent small-scale fading characteristics, such as the reduction in cluster and path counts at higher frequencies.
@@ III-B A unified channel modeling framework toward standardization
The unified channel model is illustrated in Fig. 2. The first equation describes the channel between intermediate nodes (e.g., an ISAC target or RIS panel) and the Tx/Rx. By replacing the intermediate node with a transmitter or receiver and integrating its radiation pattern and array steering vector, this equation simplifies to a direct Tx-Rx communication channel. The second equation captures the overall Tx-Target (or RIS)-Rx channel, reflecting the concatenated channel characteristics influenced by the target or RIS.
The red-boxed terms represent extensions to the 5G GBSM model, accounting for the channel characteristics of the four key 6G technologies. First, the frequency dependency of cluster number $N_{1}$ is addressed, which arises due to 6G’s potential support for wideband and new frequency communications. Without an explicit functional form for this dependency, lookup tables are used to update parameters for generating the channel’s MPCs. Unlike the fixed number of clusters in the same scenario as in the 3GPP model, the new model adjusts the cluster count based on measurement results across new mid-band to THz, as detailed in [9].
@@
Parameter $\mathrm{A}$ $(A^{Tx/Rx}_{p,n,m})$, is a path-level parameter that captures the phase differences between near-field spherical waves and far-field plane waves across the array [7]. This can be understood as the second-order term in the Taylor expansion of the array steering vector, where far-field assumptions only retain the first-order term. For model implementation, it’s essential to differentiate between LOS and NLOS components. The $\mathrm{A}$ parameter for LOS can be directly computed from the Tx/Rx geometry, while for NLOS components, intermediate scattering objects—particularly the first-bounce scatterer (FBS) and last-bounce scatterer (LBS)—must be taken into account.
@@
The fifth parameter, the Intra-Cluster K-factor, is frequency-dependent and represents the power ratio between the dominant path within a cluster and other paths. This parameter, ranging between [0,1], is derived from measurements [9]. Similar to the Rician K-factor in existing models, the Intra-Cluster K-factor allows for the redistribution of power within a cluster to simulate sparse channels. In practice, some studies model this parameter as a Gaussian-distributed statistic.
@@
The * symbol below Fig. 2 denotes a convolution operation, indicating that the Tx-Tar (or RIS)-Rx concatenated channel is the convolution of the two sub-channels in the delay domain. For each sub-channel, the angles on either side of the Target/RIS are represented using Dirichlet functions. In the concatenated channel, these Dirichlet functions must be integrated over the RCS (or radiation pattern) of the Target/RIS to determine the gain effect on signals incident from and exiting at specific angles.
The final parameter within the red box represents the RCS of the Target or the equivalent radiation pattern of the RIS. In the latest 3GPP discussions on ISAC, using RCS to model the Target has gained consensus, though the exact method for modeling RCS remains under debate. Several potential modeling approaches are under consideration: The first approach assigns fixed values based on different targets, modeling the function as an angle-independent constant, while still accounting for the four possible combinations of incident and exit polarizations. The second approach uses statistical distributions, which are more applicable to irregular objects such as UAVs, vehicles, or humans. The third method utilizes a deterministic angle-dependent function, suitable for more regular objects. For RIS, given that its application scenario emphasizes its signal modulation capabilities, modeling its behavior using a deterministic equivalent radiation pattern is both practical and effective.
III-C Implement for the unified channel model
@@
Fig. 3 shows the link-level simulation implementation of the proposed unified channel model. In particular, the top of Fig. 3 shows the current 5G standard channel model implementation process as defined by 3GPP [3], which can generally be divided into three major steps: generation of large-scale parameters such as AS and DS, generation of small-scale parameters like cluster delays, powers, angles, and intra-cluster path angles, followed by the synthesis of the channel impulse response (CIR).
@@
The implement of new model is fully built upon the framework at the bottom of the Fig. 3, and achieves the integration of different technologies by updating existing steps, adding new steps, and parameter updates. The common updates for ISAC, XL-MIMO, RIS, and new frequency communication channels are annotated within each specific implementation step.
@@
Evolution of ISAC channel. In step 1, in addition to configuring scenario layouts and Tx/Rx antenna parameters, additional configurations are required for sensing targets, especially their location, RCS characteristics, and velocity. Unlike the Tx-Rx link in traditional communication channels, ISAC systems feature a Tx-Target-Rx sensing link as well. Therefore, it is necessary to generate MPCs for the three segments: Tx-Target, Target-Rx, and Tx-Rx. These three segments exhibit certain correlated characteristics, which are introduced by the newly added step 2, generate shared scatterers ratio. According to this factor, a portion ($N_{s}$) of clusters in both communication and sensing channels are designated as shared clusters. The parameters of these clusters are then modified to maintain the similarity of MPCs across the two channels. For the sensing channel, it is also necessary to consider the concatenating of the Tx-Target and Target-Rx channels via RCS. Fig. 2 illustrates the convolution-based concatenating method. However, this method has a time complexity of $\mathrm{O}(N^{2})$, where $N$ is the number of paths in one segment of the channel. Therefore, before calculating the CIR using the convolution method (updated step 4), an additional step 3 is required to prune multipath clusters from the Tx-Target and Target-Rx channels, reducing the number of clusters and paths.
@@
Evolution of new frequency channel. The evolution of new frequency channel modeling mainly involves the extraction and update of parameters based on measurement activities. In the first step, scene-level parameters are configured. While existing 3GPP standards rely on lookup tables for generating large-scale parameters in different scenarios, the unified channel model updates these parameter tables for the new mid-band and millimeter-wave frequencies, and also adds tables for the sub-THz frequency band, based on measurement results from [9]. A significant modification concerns the number of clusters ($N_{1}$), as measurements consistently report fewer clusters than those in corresponding scenes and frequency bands in the current 3GPP model. This also impacts the generation methods for LOS probability, path loss, and other large and small-scale parameters, as outlined in [9]. In addition to modifying the parameter tables in step 1, step 3 introduces the redistribution of intra-cluster power using the ICK coefficient, which enhances the strongest path within the cluster, allowing it to occupy the majority of the cluster’s power, thereby introducing sparsity. The ICK coefficient is modeled as a Gaussian-distributed random variable based on measurement fittings, which is then used to allocate intra-cluster power. This method ultimately produces the CIR of the new frequency sparse channel through multipath combination.

### v2

# New Characteristics and Modeling of 6G Channels: Toward a Unified Channel Model for Standardization

@@ Abstract
As 6G research advances, the growing demand leads to the emergence of novel technologies such as integrated sensing and communication (ISAC), new antenna arrays like extremely large MIMO (XL-MIMO) and reconfigurable intelligent surfaces (RIS), along with multi-frequency bands (new mid-band, above 100 GHz). Standardized unified channel models are crucial for research and performance evaluation across generations of mobile communication, but the existing 5G 3GPP channel model based on geometry-based stochastic model (GBSM) requires further extension to accommodate these 6G technologies. In response to this need, this article first investigates several distinctive channel characteristics introduced by 6G technologies, such as ISAC target radar cross-section (RCS), sparsity in the new mid-band, and others. Subsequently, an extended GBSM (E-GBSM) is proposed, integrating these characteristics into a unified modeling framework. The proposed model not only accommodates 6G technologies with flexibility but also maintains backward compatibility with 5G, ensuring a smooth evolution between generations. Finally, the implementation process of the proposed model is detailed, and experiments and simulations are presented to validate its effectiveness and accuracy, providing support for 6G channel modeling standardization efforts.
@@ I Introduction
The sixth generation (6G) of wireless communication technology aspires to redefine connectivity, offering transformative advancements such as higher data rates, ultra-low latency, and ubiquitous coverage [1]. Given the complexity of 6G systems, standardization serves as a cornerstone for integrating diverse technological advancements. In particular, global standards play a pivotal role in ensuring compatibility, interoperability, and scalability across systems and regions. Notably, since 2020, the International Telecommunication Union (ITU) has spearheaded 6G research, targeting deployment by 2030, and finalized its 6G Development Framework in 2023 [2]. Concurrently, the World Radiocommunication Conference (WRC) is addressing spectrum requirements, with allocations expected by 2027. The 3rd Generation Partnership Project (3GPP) has also launched Release 19 to initiate 6G standardization efforts. As with previous generations, channel modeling remains foundational, providing the models essential for uniform benchmarking, system design, and performance evaluation.
The evolution of channel models from 1G to 6G is summarized in Table I, highlighting the transition from empirical models to standardized approaches. For 2G and earlier generations, empirical path loss models such as the Okumura-Hata model and the COST-231 model were predominant. Starting from 3G, standardized channel models have been introduced for each generation of mobile communication. The 3G era saw the introduction of the spatial channel model (SCM), combining statistical modeling with spatial correlation to support frequency-selective fading. With the widespread adoption of MIMO technology, 4G introduced the spatial channel model extension (SCME) and WINNER models, which further enhanced spatial modeling capabilities. For 5G, the geometry-based stochastic model (GBSM) integrated geometric and stochastic properties[3], while the 3D-GBSM extended channel representation into the vertical domain.
TABLE I: Development history of channel models from 1G to 6G.
@@
Dimension Time Time-frequency Time-frequency
@@
6G aims to transform communication systems by expanding application scenarios and adopting advanced technologies to meet unprecedented performance demands. The IMT-2030 Framework [2] outlines six 6G use cases, building on three primary 5G scenarios, with 15 performance indicators surpassing 5G in sensing, positioning, data rates, and coverage. Several novel technologies have been proposed to achieve these goals. Integrated sensing and communication (ISAC) combines sensing and communication functions for precise positioning and autonomous driving. Advanced antenna technologies, such as extremely large MIMO (XL-MIMO) and reconfigurable intelligent surfaces (RIS), enhance spectral efficiency, capacity, and reliability. XL-MIMO utilizes extensive antenna arrays, while RIS actively manipulates electromagnetic waves to improve coverage. To support these capabilities, new frequency ranges, such as the new mid-band, provide the bandwidth and spectrum resources required to achieve 6G’s peak data rates [4].
While the 5G channel model, grounded in GBSM, has successfully supported key technologies like MIMO and millimeter-wave communication, it now cannot be directly applied to 6G due to expanded technical requirements, broader frequency ranges, and more diverse scenarios introduced by 6G technologies. ISAC systems, for instance, necessitate target modeling to capture the characteristics of both communication and sensing channels[5]. XL-MIMO’s large antenna arrays introduce spatial non-stationarity (SnS) and near-field effects[6], significantly altering channel behavior. RIS-assisted communications present challenges in modeling the transmitter (Tx)-RIS and RIS-receiver (Rx) channel coupling, while new frequency bands, such as mid-band [7], millimeter-wave, and sub-terahertz (sub-THz), exhibit sparse multipath propagation[8].
@@
In response to these challenges, this article investigates the channel characteristics introduced by key 6G technologies and, drawing on discussions from the 3GPP Technical Specification Group Radio Access Network Working Group 1 (TSG RAN1), proposes a channel modeling framework based on an extended GBSM (E-GBSM). This framework moves toward a unified modeling approach by integrating the distinctive features of ISAC, advanced antenna technologies, and new frequency bands (new mid-band and above 100 GHz), while ensuring compatibility with existing standards. By extending current models, it maintains backward compatibility, provides a clear implementation roadmap, and addresses critical requirements such as wideband support and multi-technology evaluation.
@@ II-A ISAC
Compared to traditional communication channels, ISAC channels emphasize the behavior and characteristics of sensing targets. Measurement studies focus on three aspects: (1) radar cross-section (RCS) properties of sensing targets, (2) modeling of the target channel and its sub-channels (Tx-target and target-Rx), and (3) correlations between sensing target channels and communication channels.
The first aspect involves the investigation of RCS properties. Although extensively studied in radar systems, ISAC targets differ in frequency bands and scenarios, necessitating further research [10]. Typical targets include unmanned aerial vehicle (UAV), human, and vehicle, relevant to applications such as UAV detection and autonomous driving assistance. Measurements reveal distinct RCS distributions, e.g., small UAVs and humans exhibit weak angular dependence, while vehicles show four peaks corresponding to their sides. RCS also varies with frequency and distance, stabilizing at longer ranges. Certain studies highlight potential correlations between bistatic and monostatic RCS within specific angular ranges, simplifying modeling.
@@ II-B New antenna
The channel characteristics of advanced antenna technologies, such as XL-MIMO and RIS, focus on near-field effects, SnS, and multi-segment concatenated channels in RIS systems.
@@ II-C New frequency
As carrier frequency increases and the signal wavelength decreases, two propagation effects emerge. First, the diffraction capability of shorter wavelengths diminishes, resulting in a stronger reliance on line-of-sight (LOS) and a limited number of reflected paths. This reduces the number of resolvable clusters in the environment. Second, the smaller wavelength enlarges the relative scale of surface irregularities on objects, leading to more angularly dispersed and lower-power scattering components. As a result, the intra-cluster power tends to concentrate in a few dominant reflected paths. Despite these frequency-dependent characteristics, current 3GPP channel models do not reflect such sparsity. Across a wide frequency range—from 0.5 to 100 GHz—the models assume a fixed number of clusters for a given scenario and assign equal power to 20 paths within each cluster. Various metrics, such as channel rank or spatial degrees of freedom, could quantify sparsity. This article adopts the Gini Index, as proposed in [8], to measure sparsity, where a Gini Index closer to 1 (within the 0-1 range) indicates greater sparsity. Channel measurements from [8] across centimeter-wave, new mid-band, and THz bands show that the Gini Index increases with frequency, confirming that higher frequencies exhibit more pronounced sparsity.
Figure 2: The illustration of the E-GBSM for 6G.
@@ III-A Consideration for 6G channel model
Support for evaluating multiple technologies. While the 5G channel model accommodates several features like spatial consistency, temporal evolution, and blockage effects (e.g., vehicle-to-everything (V2X) networks and oxygen absorption for 53-67 GHz), it lacks provisions for many 6G technologies. For instance, ISAC requires models that capture sensing channels from the Tx to the target and then to the Rx, which are unsupported in current models like 3GPP TR 38.901 [3]. Similarly, RIS models must represent concatenated Tx-RIS-Rx channels, and XL-MIMO demands adjustments to incorporate non-stationary and near-field effects, which traditional models based on spatial stationarity and far-field assumptions fail to address.
Support for a wide frequency range. Achieving 6G’s peak data rates requires both higher frequencies (e.g., THz bands [11]) and better utilization of underused bands (e.g., new mid-band). Current models extrapolate new mid-band parameters from FR1 and FR2 bands [12] and lack representation of the frequencies above 100 GHz. Moreover, they do not account for frequency-dependent small-scale fading characteristics, such as the reduction in cluster and path counts at higher frequencies.
@@ III-B A unified channel modeling framework toward standardization
The unified channel model is illustrated in Fig. 2. The first equation characterizes the channel between the $p$th transmit (Tx) antenna and the $x$th intermediate node (either an ISAC target or a RIS panel). Here, $\mathbf{F}_{p}$ is a $2\times 1$ vector, and $\mathbf{CPM}_{n_{t},m_{t}}$ is a $2\times 2$ matrix. The resulting $2\times 1$ vector $\mathbf{r}_{p,x}^{\mathrm{Tx{-}Tar}}$ represents the angle-domain and delay-domain channel impulse response (CIR) for the two polarization components. Similarly, for the Target–Rx link, denoted as $\mathbf{r}_{q,x}^{\mathrm{Rx{-}Tar}}(\Omega^{\mathrm{out}},\tau)$, the expression follows the same structure as in the Tx–Target case, with the following substitutions: Tx $\rightarrow$ Rx, $p\rightarrow q$, $n_{t}\rightarrow n_{r}$, and $m_{t}\rightarrow m_{r}$. The second equation captures the overall Tx-target (or RIS)-Rx channel, reflecting the concatenated channel characteristics influenced by the target or RIS.
The red-boxed terms represent extensions to the 5G GBSM model, accounting for the channel characteristics of the four key 6G technologies. First, the frequency dependency of cluster number $N_{1}$ is addressed, which arises due to 6G’s potential support for wideband and new frequency communications. Without an explicit functional form for this dependency, lookup tables are used to update parameters for generating the channel’s multipath components (MPCs). Unlike the fixed number of clusters in the same scenario as in the 3GPP model, the new model adjusts the cluster count based on measurement results across new mid-band to THz, as detailed in [8].
@@
Parameter $\mathrm{A}$ $(A^{Tx/Rx}_{p,n,m})$, is a path-level parameter that captures the phase differences between near-field spherical waves and far-field plane waves across the array [6]. This can be understood as the second-order term in the Taylor expansion of the array steering vector, where far-field assumptions only retain the first-order term. For model implementation, it’s essential to differentiate between LOS and non-line-of-sight (NLOS) components. The $\mathrm{A}$ parameter for LOS can be directly computed from the Tx/Rx geometry, while for NLOS components, intermediate scattering objects—particularly the first-bounce scatterer (FBS) and last-bounce scatterer (LBS)—must be taken into account.
@@
The fifth parameter, the intra-cluster K-factor (ICK), is frequency-dependent and represents the power ratio between the dominant path within a cluster and other paths. This parameter, ranging between [0,1], is derived from measurements [8]. Similar to the Rician K-factor in existing models, the ICK allows for the redistribution of power within a cluster to simulate sparse channels. In practice, some studies model this parameter as a Gaussian-distributed statistic.
@@
The * symbol below Fig. 2 denotes a convolution operation, indicating that the Tx-target (or RIS)-Rx concatenated channel is the convolution of the two sub-channels in the delay domain. For each sub-channel, the angles on either side of the target/RIS are represented using Dirichlet functions. In the concatenated channel, these Dirichlet functions must be integrated over the RCS (or radiation pattern) of the target/RIS to determine the gain effect on signals incident from and exiting at specific angles.
The parameter $F^{\mathrm{Tar}}$ represents either the square root of the target’s RCS or the equivalent radiation pattern of a RIS. Recent 3GPP discussions on ISAC have converged to a consensus on the specific parametrization of monostatic RCS, which is formulated as: $F^{\mathrm{Tar}}=A\cdot B_{1}(\Omega^{\mathrm{out}},\Omega^{\mathrm{in}})\cdot B_{2},$ where $A$ denotes the total reflected power intensity of the target, $B_{1}$ characterizes the angle-dependent scattering component, and $B_{2}$ represents a statistical distribution term modeling the random fluctuation of RCS. The term $\mathrm{CPM}^{\mathrm{Tar}}$ describes the polarimetric interaction between the target and the incident signal. Its mathematical structure aligns with the polarization matrix defined in the 3GPP TR 38.901 standard, adopting a $2\times 2$ Jones matrix formalism.
III-C Implementation for the unified channel model
@@
Fig. 3 shows the link-level simulation implementation of the proposed unified channel model. In particular, the top of Fig. 3 shows the current 5G standard channel model implementation process as defined by 3GPP [3], which can generally be divided into three major steps: generation of large-scale parameters such as AS and DS, generation of small-scale parameters like cluster delays, powers, angles, and intra-cluster path angles, followed by the synthesis of the CIR.
@@
The implementation of new model is fully built upon the framework at the bottom of the Fig. 3, and achieves the integration of different technologies by updating existing steps, adding new steps, and modifying parameters. The common updates for ISAC, XL-MIMO, RIS, and new frequency communication channels are annotated within each specific implementation step.
@@
Evolution of ISAC channel. In step 1, in addition to configuring the scenario layout and Tx/Rx antenna parameters, further settings are needed for sensing targets, particularly their location, RCS, and velocity. Unlike conventional communication channels that involve only a Tx-Rx link, ISAC systems also include a Tx-target-Rx sensing link. Thus, MPCs must be generated for three segments: Tx-target, target-Rx, and Tx-Rx. These segments exhibit correlated characteristics, which are addressed in step 2 through the generation of shared scatterers. A portion ($N_{s}$) of clusters in both the communication and sensing channels are designated as shared, with their parameters adjusted to preserve MPC similarity. Compared with the authors’ prior work, which primarily focused on modeling shared scatterers between communication and sensing channels, the ISAC component in Fig.3 highlights a distinct feature: the modeling of the sensing channel as the concatenation of the Tx-target and target-Rx channels via RCS. Fig.2 illustrates the proposed convolution-based concatenation method. However, this approach has a computational complexity of $\mathcal{O}(N^{2})$, where $N$ denotes the number of paths in one segment of the channel. To address this, an additional step—step 3—is introduced before computing the CIR in updated step 4. This step prunes excess multipath clusters from the Tx-target and target-Rx segments, thereby reducing the number of clusters and paths involved.
@@
Evolution of new frequency channel. The modeling of new frequency channels primarily involves updating parameters based on measurement activities. First, the scene-level parameters are configured. While existing 3GPP standards use lookup tables to generate large-scale parameters for different scenarios, the unified channel model updates these tables for the new mid-band and millimeter-wave frequencies, and adds entries for the sub-THz band based on measurements reported in [8]. A significant modification concerns the number of clusters ($N_{1}$), as measurements consistently report fewer clusters than those in corresponding scenes and frequency bands in the current 3GPP model. This also impacts the generation methods for LOS probability, path loss, and other large and small-scale parameters, as outlined in [8]. Beyond updating parameter tables, step 3 introduces intra-cluster power redistribution via the ICK, enhancing the strongest path within a cluster to dominate the power and thus introducing sparsity. The ICK coefficient is modeled as a Gaussian-distributed random variable based on measurement fittings, which is then used to allocate intra-cluster power. This method ultimately produces the CIR of the new frequency sparse channel through multipath combination.